OpenStudy (anonymous):

is f(x)=xabs(x^2) an odd, even or neither function

OpenStudy (anonymous):

even i belive

OpenStudy (anonymous):

you just look at the expontes

OpenStudy (anonymous):

it is online hw and i only have one attempt to get it right

OpenStudy (anonymous):

u have to be 100% sure

OpenStudy (anonymous):

\[f(x)=x \left| x^2 \right|\]

OpenStudy (anonymous):

is it odd, even or neither

OpenStudy (anonymous):

So in theory, IF: \(f(x)=f(-x)\) THEN It is even. \(-f(x)=f(-x)\) THEN it is odd.

OpenStudy (anonymous):

So if we know that \(f(x)=x|x^2|\), Lets check to see if the identities are correct! Try computing f(-x) and -f(x) and see what those return and if it fits any of the above equations.

OpenStudy (anonymous):

can u give me an answer?

OpenStudy (anonymous):

please

OpenStudy (anonymous):

i get confused with the absolute value

OpenStudy (anonymous):

I don't get what hes saying what my teacher told me is you just look at the exponets

OpenStudy (anonymous):

Oh yeah...good point. Try this: \[f(x)=x|x^2|\] \[f(-x)=(-x)|(-x)^2|=-x|x^2|\] \[-f(x)=-(x)|x^2|=-x|x^2|\] So since \(f(-x)=-f(x)\) The function is odd.

OpenStudy (anonymous):

I am \(200\%\) sure that \(f(x)=x|x^2|\) is odd

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

what do i do if it gives me two graphs

OpenStudy (anonymous):

If what does...?

OpenStudy (anonymous):

the hw

OpenStudy (anonymous):

don't know if the link works

OpenStudy (anonymous):

@KeithAfasCalcLover

OpenStudy (anonymous):

good luck t(-_-)t

OpenStudy (anonymous):

i wil do it on my own i just want u to help me start

OpenStudy (anonymous):

And what you want to do with the graph?

OpenStudy (anonymous):

find out if they r odd, even or neither

OpenStudy (anonymous):

@KeithAfasCalcLover

OpenStudy (anonymous):

Visually you can identify a function as odd even or neither by imagining this: If the right or left side of \(f(x)\) can be rotated about the y-axis to match the other side, it is even. If the If the right or left side of \(f(x)\) can be rotated about the line of \(y=x\) or \(y=-x\) to match the other side, it is odd. Otherwise it is neither.

OpenStudy (anonymous):

so what do u think

OpenStudy (anonymous):

lol why don't you think instead of just asking for answers

OpenStudy (anonymous):

i think its red=even and blue=odd. what do u think

OpenStudy (anonymous):

Well think about it, could you flip one side of the red over the y axis to match the other?

OpenStudy (anonymous):

have u looked at the graph

OpenStudy (anonymous):

Yes muzzamil, I have.

OpenStudy (anonymous):

buddy i really need help on this

OpenStudy (anonymous):

LOL KEITH WALK AWAY NOW

OpenStudy (anonymous):

Have you? Take a look. Imagine you grab one side and fold it over the y-axis. Does it fit on the other side?

OpenStudy (anonymous):

i said i think red is even and blue is odd. atlesat i am trying

OpenStudy (anonymous):

i asked if u thouht i was right

OpenStudy (anonymous):

Well why do you think that?

OpenStudy (anonymous):

and he is asking you a question and you haven't answered it and if you did it answers your question llol

OpenStudy (anonymous):

kramer instead of wasting ur precious time maybe u should leave. i did what u told me to

OpenStudy (anonymous):

trust me my time at the moment is not precious

OpenStudy (anonymous):

i keep asking ur opinion cuz my a$$ is on the line. my parents get pissed if i mess up on my hw. its an asian thing

OpenStudy (anonymous):

didn't you get taught the lesson in school?

OpenStudy (anonymous):

i also want to watch football this weekend and i need this done. i am a ninth grader taking calculus at the university of minnesota. the professors suck there

OpenStudy (anonymous):

bro that isn't calculus that's pre calc

OpenStudy (anonymous):

its th first chapter of the book

OpenStudy (anonymous):

u dont do derivatives the first day

OpenStudy (tkhunny):

x is odd. x^2 is even abs(x^2) is redundant because x^2 is always positive. Therefore, f(x) = x^3 and it is definitely odd.

OpenStudy (anonymous):

we already solved that. we r on the graph one. i posted a link @tkhunny

OpenStudy (anonymous):

I would have to agree that it is not calculus but it may be review of pre calc in the beginning. How youre accepted in Minnesota University at grade 9 is beyond me.

OpenStudy (anonymous):

because hes a liar lol

OpenStudy (anonymous):

UMTYMP. university of minnesota talented youth math program

OpenStudy (anonymous):

look it up. mathcep.umn.edu

OpenStudy (anonymous):

@llkramer mathcep.umn.edu

OpenStudy (tkhunny):

I saw the solution. I just wanted to make sure we used the word "redundant".

OpenStudy (anonymous):

@tkhunny this guy just wants answers no work don't waste your time

OpenStudy (anonymous):

@llkramer did u look it up?

OpenStudy (anonymous):

\[f(x)=x|x|^2\] \[f(2)=2\times |2|^2=\] \[f(-2)=-2\times |-2|^2=-8\]

OpenStudy (anonymous):

we already solved this @satellite73 . we r now on the link i posted with the two graphs

OpenStudy (anonymous):

DONT GIVE HIM A ANSWER BRO

OpenStudy (anonymous):

what a pretty picture what are you supposed to do with it?

OpenStudy (anonymous):

AN**

OpenStudy (anonymous):

r the graphs even, odd, or neither

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

i think red=even and blue=odd

OpenStudy (anonymous):

the one that is symmetric with respect to the \(y\) axis is even the one that is symmetric with respect to the origin \((0,0)\) is odd

OpenStudy (anonymous):

yes, you are right

OpenStudy (anonymous):

LOL @satellite73 "What a pretty picture" haha that got a good laugh a out of me.

OpenStudy (anonymous):

THANK YOU @satellite73 FINALLY SOMEBODY WHO COULD ACTUALLY TELL ME IF I WAS RIGHT OR WRONG.

OpenStudy (anonymous):

@satellite73 That's EXACTLY what I thought when It was first shown to me

OpenStudy (anonymous):

THEN Y COULDN''T U SAY THAT

OpenStudy (anonymous):

U COULD HAVE SAVED US LIKE 30 MIN

OpenStudy (anonymous):

Though math is about right and wrong, It is even more about the WHY. "yes" and "no" wont get you very far without the WHY.

OpenStudy (anonymous):

WELL THANKS FOR THE FIRST ONE ANYWAY

OpenStudy (anonymous):

Is this some sort of an "ALL CAPS RAGE" or so?

OpenStudy (anonymous):

idk. bye

OpenStudy (anonymous):

Bye. GOOD LUCK